OpenStudy (anonymous):

solve for x \[8^{x-1} = 2^{x+1}\]

OpenStudy (phi):

replace 8 with 2^3 and go from there...

OpenStudy (anonymous):

0.o?

OpenStudy (anonymous):

\[2^{3^{x-1}} = 2^{x+1}\] ?

OpenStudy (anonymous):

rule of exponents i mean

OpenStudy (phi):

Pretty close. 3*(x-1) not 3x-1

OpenStudy (anonymous):

it's not 3^x-1?

OpenStudy (phi):

\[(2^{3})^{x-1}= 2^{3(x-1)}\]

OpenStudy (anonymous):

wait I seem to remember if the bases are the same ( both 2) then it should just be 3(x-1) = x+1 right?

OpenStudy (phi):

yes.

OpenStudy (anonymous):

x=2?

OpenStudy (phi):

Yes. But you can (and should) check it. plug 2 into the original problem

OpenStudy (anonymous):

gives me 8 = 2 :(

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

hey it's right!

OpenStudy (anonymous):

thanks phi!